OpenStudy (anonymous):

The center of mass of a 0.2 kg (non-uniform) meter stick is located at its 46-cm mark. What is the magnitude of the torque (in Nm) due to gravity if it is supported at the 26-cm mark? (Use g = 9.79 m/s2).

OpenStudy (4n1m0s1ty):

So here's your stick: ---------26cm----------------------------46cm------------------- fulcrum center of mass where are the forces acting on this system?

OpenStudy (anonymous):

Torque=Force*Distance Weight=m*g Distance=46-26 Force=?

OpenStudy (anonymous):

Remember to put the distance in m, so distance=0.2m